OpenStudy (anonymous):

Choose the slope-intercept: equation of the line that passes through the point (6, -6) and is perpendicular to y = 3x - 6. A) y = 1/3x - 8 B) y = -3x + 12 C) y = 3x - 24 D) y = -1/3 x - 4

OpenStudy (anonymous):

slope intercept form ; y=mx+b here slope for the line y=3x-6 is 3 but since we want the slope of a line perpendicular to this::::: so negative recip[rocal of this is taken...... so here m=-1/3

OpenStudy (anonymous):

right till here...........?

OpenStudy (anonymous):

y=mx+b -6=-1/3(6)+b -6=-2+b -4=b b=-4 so the equation is: y=(-1/3)x-4

OpenStudy (anonymous):

@chariety101

OpenStudy (anonymous):

oh I'm sorry i wasn't responding, I actually did get it a lot better. Thanks so much!